OpenStudy (anonymous):

Find the product:(x^2-2y)(x^2+2y)

OpenStudy (saifoo.khan):

\[x^4-4 y^2 \]

OpenStudy (anonymous):

since \[(a+b)(a-b)=a^2-b^2\] therefore \[(x^2-2y)(x^2+2y)=(x^2)^2-(2y)^2\] \[x^4-4y^2\]